a rectangle has a length of 6 cm. what is its width if its perimeter and area are the same number?
Answer by checkley71(8403)

6W=2*W+2*6
6W=2W+12
6W-2W=12
4W=12
W=12/4
W=3 ANSWER FOR THE WIDTH.
PROOF:
6*3=2*3+2*6
18=6+12
18=18